40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 14 13:29:51 2018, gautam, Update, General, PSL shutter closed, all optics misaligned 
    Reply  Thu Jun 14 19:30:12 2018, gautam, Update, General, IFO alignment restored 
       Reply  Thu Jun 14 22:45:05 2018, johannes, Update, General, AUX beam SRC alignment 
          Reply  Sat Jun 16 00:26:48 2018, gautam, Update, General, PRC modescan attempt 
             Reply  Sat Jun 16 01:25:29 2018, Koji, Update, General, PRC modescan attempt 
                Reply  Sat Jun 16 20:57:59 2018, Jon, Update, General, PRC modescan attempt 40M_SRMI_AM_annotated.pdf40M_DRMI_AM.pdf
Message ID: 13962     Entry time: Thu Jun 14 13:29:51 2018     Reply to this: 13967
Author: gautam 
Type: Update 
Category: General 
Subject: PSL shutter closed, all optics misaligned 

[jon, gautam]

Jon is doing some characterization of the AUX laser setup for which he wanted only the prompt retroreflection from the SRM on the AS table, so the PSL shutter is closed, and both ITMs and ETMs are misaligned. The prompt reflection from the SRM was getting clipped on something in vacuum - the ingoing beam looked pretty clean, but the reflection was totally clipped, as I think Johannes aligned the input beam with the SRM misaligned. So the input steering of the AUX laser beam into the vacuum, and also the steering onto AS110, were touched... Also, there were all manner of stray, undumped beams from the fiber on the AS table noJon will post photos.

Before we began this work, we found that c1susaux was dead so we rebooted it.

ELOG V3.1.3-